### v1

# CaSTFormer: Causal Spatio‐Temporal Transformer for Driving Intention Prediction

Abstract
Accurate prediction of driving intention is key to enhancing the safety and interactive efficiency of human-machine co-driving systems. It serves as a cornerstone for achieving high-level autonomous driving. However, current approaches remain inadequate for accurately modeling the complex spatio-temporal interdependencies and the unpredictable variability of human driving behavior. To address these challenges, we propose CaSTFormer, a Causal Spatio-Temporal Transformer to explicitly model causal interactions between driver behavior and environmental context for robust intention prediction. Specifically, CaSTFormer introduces a novel Reciprocal Shift Fusion (RSF) mechanism for precise temporal alignment of internal and external feature streams, a Causal Pattern Extraction (CPE) module that systematically eliminates spurious correlations to reveal authentic causal dependencies, and an innovative Feature Synthesis Network (FSN) that adaptively synthesizes these purified representations into coherent spatio-temporal inferences. We evaluate the proposed CaSTFormer on the public Brain4Cars dataset, and it achieves state-of-the-art performance. It effectively captures complex causal spatio-temporal dependencies and enhances both the accuracy and transparency of driving intention prediction.
Introduction
Driver intention prediction is critical in autonomous driving systems, playing a significant role in effectively mitigating risks and enhancing driving safety. By forecasting potential outcomes several seconds in advance, the system can alert the driver proactively or initiate evasive maneuvers, further increasing its safety capabilities.
Typically, the intention prediction relied on the extraction and fusion of visual features to perform simple intention prediction (Huang et al. 2022). The development of sensor technology has enabled the use of multi-modal information, including GPS coordinates, vehicle speed, map data, and driver head pose (Hu et al. 2021; Mo et al. 2023; Li, Zhao, and Wang 2022; Wu et al. 2023). By leveraging complex models for feature extraction, fusion, and prediction, the performance of driver intention prediction has been significantly improved (Sui et al. 2021; Guo et al. 2023; Liu, Wu, and Wang 2023; Gao et al. 2023).
Figure 1: Comparison between previous driving intention prediction methods and ours. (a) is an LSTM–GRU–based framework in which internal and external streams are processed independently and then concatenated for final fusion. (b) is our CaSTFormer, a Transformer–based model that performs dynamic causal fusion of the dual streams while integrating intention priors. By jointly capturing global and local dependencies and leveraging intrinsic cross-stream interactions, our method surpasses existing approaches.
Despite the progress made in driver intention prediction, the rich multi-modal data has not been fully exploited. Most existing work merely concatenates or linearly aggregates this information, as illustrated in Fig. 1 (a). However, as the controller of the vehicle, changes in the driver’s state subsequently influence the vehicle’s driving status, revealing a tight dependency. Therefore, we propose to explicitly model the causal relationship between the driver and the environment, emphasizing the decisive impact of this causality on the prediction task, as shown in Fig 1 (b).
Specifically, we adopt a transformer-based architecture and introduce three sequential components to enhance the causal modeling, encode the driver’s intention, and fuse the multi-dimensional features. First, we introduce a Reciprocal Shift Fusion (RSF) module that cross-fuses internal and external features through a shifting mechanism, explicitly establishing a temporal dependency between the two feature streams. As the fused features may contain considerable casual noise, we further devise a Causal Pattern Extraction (CPE) module to filter out such noise to obtain a clearer causal representation. Finally, we utilize a Feature Synthesis Network (FSN) that employs a gating mechanism to integrate the in-cabin, out-of-cabin, and interaction representations, enabling the modeling of both local and global causal structures. Our main contributions are as follows:
  • We propose CaSTFormer, an efficient Transformer-based framework for driving-intent prediction that integrates temporal dependency modeling, causal pattern enrichment, and adaptive multi-view fusion within a unified end-to-end architecture.
  • CaSTFormer disentangles spatio‐temporal interactions of interior and exterior streams, subsequently isolates genuine causal contributions through counterfactual attention subtraction and adaptively fuses heterogeneous features, thereby mitigating spurious correlations and improving robustness under complex driving conditions.
  • Extensive evaluation on the Brain4Cars dataset demonstrates that CaSTFormer demonstrates superior performance in driving intention prediction in both highway and urban scenarios.
Related Work
In the early stages of research on driving intention prediction, some studies focused on learning spatio-temporal features directly from raw video data. For instance, a combination of 3D Convolutional Neural Networks and Long Short-Term Memory networks was utilized to predict driving maneuvers directly from video streams (Gebert et al. 2019). To achieve a more comprehensive basis for decision-making, multi-modal data was introduced to improve the accuracy of intention prediction. In one such study, the in-cabin driver’s view was combined with the external traffic scene, using a Convolutional Long Short-Term Memory network to fuse and analyze features from both views (Rong, Akata, and Kasneci 2020). However, their contextual capacity is limited, making it difficult to process long-range, non-consecutive events.
Inspired by human cognitive processes, TIFN (Guo et al. 2023) proposed a State Update Unit to incorporate the influence of environmental information into driver state modeling, and extracted semantic segmentation features to provide clear cues affecting driver attention. Meanwhile, another approach treated intention prediction as a sequence labeling task, further capturing the contextual dependencies of driving behavior by combining a Bidirectional Long Short-Term Memory network with a Conditional Random Field (Zhou et al. 2021). Although these models fuse multi-source information, they typically learn the correlations between features implicitly, without explicitly decoupling and reasoning about the mutual influences between different information modalities.
Furthermore, personalized differences among drivers were also taken into consideration as a significant source of model error. Early work used a Domain-Adversarial Recurrent Neural Network (DA-RNN) to learn driver-agnostic, domain-invariant features to enhance the model’s generalization capabilities (Tonutti et al. 2019). Recently, IDIPN (Liu et al. 2025) utilized Inverse Reinforcement Learning to extract unique driving preferences from each driver’s historical data, thereby correcting the predictions of a general-purpose model, while FedPRM (Zhu et al. 2024) achieved this by using a federated learning framework to directly construct personalized models for each driver.
The Transformer architecture, with its unique attention mechanism, excels at capturing long-range dependencies and global interaction information in temporal data (Liu et al. 2024). CemFormer (Ma et al. 2023) integrates data from both in-cabin and external cameras, learning a unified cross-view representation via a spatio-temporal tansformer to infer driver intention directly from their behavior. However, such methods primarily rely on the inherent structure of the transformer to interpret the dependencies. Meanwhile, a non-autoregressive transformer with hybrid attention has been utilized to simultaneously capture the temporal dynamics of a single vehicle, as well as the interactions among multiple vehicles (Jiang et al. 2024). DriveTransformer (Jia et al. 2025) further constructs a unified end-to-end framework to handle perception, prediction, and planning tasks in parallel. This treats intention prediction as an integrated component of the overall scalable system, but may dilute the model’s focus on driver-intention features.
Besides, some early works have attempted to introduce causal inference methods into the intention prediction task. One study constructed a causal model to learn the temporal relationships of invariant representations from driving data, thereby achieving domain generalization (Hu et al. 2022). Another adopted a driver-centric approach, formulating the task of risk object identification as a causal problem and proposing a two-stage framework based on causal inference (Li, Chan, and Chen 2020). However, the scope of scenarios and conditions covered by these methods is relatively limited.
Apart from all the prior work, our method employs a dual-stream transformer. Building upon an encoding of driver intention, it explicitly establishes both local and global causal dependencies between in-cabin and external factors. This approach, through the fusion of multi-dimensional features, enables precise intention prediction to address increasingly complex driving scenarios.
Figure 2: Overview of the CaSTFormer pipeline. After data preprocessing, external optical flow is encoded by ResNet-18 and interior images by MobileFaceNet to produce dual-stream feature sequences. These are then fed into three core modules: (a) Reciprocal Shift Fusion (RSF) for temporal feature integration; (b) Causal Pattern Extraction (CPE) for causal enhancement and intent embedding; and (c) Feature Synthesis Network (FSN) for dynamic fusion of complementary internal, external and interaction views to yield the final driving intention prediction.
Method
As illustrated in Fig. 2, CaSTFormer takes as input a bi-stream image sequence =$\mathcal{I}=\{(I^{\mathrm{out}}_{b,t},I^{\mathrm{in}}_{b,t})\}_{b,t=1}^{B,T}$ comprising $B$ samples, each consisting of $T$ frames from two synchronized streams (outside and inside views), where $b$ and $t$ index the sample and temporal frame, respectively. Feature extractors separately encode each stream into features $F^{\mathrm{out}},F^{\mathrm{in}}\in\mathbb{R}^{B\times T\times D}$ ($D$ denotes the dimensionality of each encoded feature vector). The obtained features pass through a Reciprocal Shift Fusion (RSF) module for temporal causality modeling via our proposed bi-stream attention, followed by a Causal Pattern Extraction (CPE) module to inject learnable causal representations. Finally, visual features ($z_{\mathrm{in}},z_{\mathrm{out}}$ and interaction features $z_{\mathrm{ctx}}$) are adaptively fused by a Feature Synthesis Network (FSN) into a joint prediction $\ell_{\mathrm{joint}}$.
Reciprocal Shift Fusion (RSF)
On multi‐lane highways, cameras concurrently record the external traffic scene and the driver’s internal state. Recognizing the mutual influences between environmental context and driver behavior, our dual‐stream architecture explicitly captures their bidirectional interactions by jointly modeling both feature streams. To enforce exterior-first precedence, we introduce a temporal delay mechanism in the Key and Value sequences. Specifically, at time step $t$, the attention mechanism accesses only information from the preceding frame $t-1$. Concretely, we define the delayed feature
[…]
applied separately to $F^{\mathrm{out}}$ and $F^{\mathrm{in}}$.
Bidirectional Dependency Attention (BDA).
Under a strict single-frame delay constraint, BDA enriches each frame’s representation by fusing internal and external contexts from the immediately preceding timestep. The current internal and external features attend bidirectionally to their one-frame–delayed counterparts, capturing both temporal coherence and cross-stream coupling. To model diverse associations efficiently, we project into $H$ parallel attention heads and aggregate their outputs through concatenation and a final linear mapping:
$\mathrm{BDA}(Q,K,V)=\Bigl{[}\mathrm{softmax}\bigl{(}\tfrac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}\bigr{)}V_{i}\Bigr{]}_{i=1}^{H}W^{O},$ (2)
where $[\cdot]_{i=1}^{H}$ denotes head-wise concatenation and $W^{O}$ restores the original feature dimension. Figure 3 illustrates the detailed bidirectional query–key–value fusion, showing how the internal and external streams are jointly refined.
Figure 3: Illustration of Bidirectional Dependency Attention (BDA) where buffered internal and external features are mutually cross-attended to produce dynamically enhanced current-frame representations.
Channel-wise Gating.
Although BDA produces a richly fused representation $H_{b,t}$ at each spatial–temporal location, individual channels may still carry irrelevant or noisy information. Hence, we employ two successive channel‐wise gating layers that adaptively emphasize informative channels and suppress spurious responses:
$g_{b,t}$ $=\sigma\bigl{(}W_{2}\,(\mathrm{ReLU}(W_{1}H_{b,t}+b_{1}))+b_{2}\bigr{)}$ (3)
$\widetilde{H}_{b,t}$ $=g_{b,t}\,\odot\,H_{b,t}$
where $\odot$ denotes the Hadamard product between vectors.
Normalization and Regularization.
For numerical stability and to guard against overfitting, the following operations are applied to the gated outputs $\widetilde{H}_{b,t}$:
$R(x)=x\;\odot\;\frac{s}{\sqrt{\frac{1}{D}\sum_{d=1}^{D}x_{d}^{2}+\epsilon}}$ (4)
${X}_{b,t}=\mathrm{Dropout}\bigl{(}R(\widetilde{H}_{b,t})\bigr{)}$ (5)
where $R(\cdot)$ denotes the Root-mean-square normalization function, $s$ is a learnable scaling parameter, $\epsilon$ is a small constant to ensure numerical stability, and $D$ is the dimensionality of the feature vector $x$, yielding channel-calibrated representations across both internal and external feature streams.
Causal Pattern Extraction (CPE)
Conventional intention prediction architectures aggregate heterogeneous interior and exterior cues under an implicit correlation assumption and thus often mistake coincidental patterns for true decision drivers. To overcome this limitation, our CPE module contrasts observed and counterfactual cross‐stream attentions to disentangle direct causal contributions, then selectively amplifies only those residuals that genuinely influence driving intent, resulting in improved robustness and generalization in safety‐critical scenarios. Specifically, CPE takes the bidirectionally fused interior and exterior features ${X}_{b,T}^{\mathrm{in}},{X}_{b,T}^{\mathrm{out}}\in\mathbb{R}^{T\times B\times D}$ as inputs to perform this causal reasoning process.
Direct Causal Effect.
At each time step, we compute two attention distributions. We first calculate observed attention $A^{\mathrm{obs}}_{\mathrm{in},t}$ using actual exterior features, and then generate counterfactual attention $A^{\mathrm{cf}}_{\mathrm{in},t}$ by replacing all exterior features with their temporal mean $\bar{X}^{\mathrm{out}}=\frac{1}{TB}\sum_{u=1}^{T}\sum_{b=1}^{B}X^{\mathrm{out}}_{u,b}$, which serves as a neutral baseline that removes environmental variations. The difference $\Delta^{\mathrm{in}}_{t}$ between these two distributions quantifies the direct causal influence of external context on interior representations. Specifically, for $t=1,\dots,T$, we obtain
$A^{\mathrm{obs}}_{\mathrm{in},t}$ $=\mathcal{A}\bigl{(}X^{\mathrm{in}}_{t},\,X^{\mathrm{out}}_{\leq t-1},\,X^{\mathrm{out}}_{\leq t-1}\bigr{)},$ (6)
$A^{\mathrm{cf}}_{\mathrm{in},t}$ $=\mathcal{A}\bigl{(}X^{\mathrm{in}}_{t},\,\bar{X}^{\mathrm{out}},\,\bar{X}^{\mathrm{out}}\bigr{)},$
$\Delta^{\mathrm{in}}_{t}$ $=A^{\mathrm{obs}}_{\mathrm{in},t}-A^{\mathrm{cf}}_{\mathrm{in},t},$
where $\mathcal{A}(Q,K,V)$ denotes multi‐head scaled dot‐product attention. Similarly, $\Delta^{\mathrm{out}}_{t}$ is defined by interchanging the two streams. We further orthogonalize each causal residual against the global baseline vector $\bar{X}$ to ensure that the identified causal patterns reflect true intent-relevant dependencies rather than dataset-specific biases. Formally,
$\Delta_{t}^{\perp}=\Delta_{t}-\frac{\langle\Delta_{t},\,\bar{X}\rangle}{\|\bar{X}\|^{2}+\varepsilon}\,\cdot\bar{X},$ (7)
where $\varepsilon$ is a small constant to ensure numerical stability. The orthogonal projection produces $\Delta^{\perp}_{t}$ that captures context-deviating residuals while removing baseline-aligned components. We retain the final-step residual $\Delta^{\perp}_{T}$ as the decision-relevant signal for downstream modules, ensuring that subsequent fusion operates exclusively on temporally salient causal information.
Dynamic Causal Gating.
The causal relevance of residuals differs across driving scenarios, as critical maneuvers merit amplification, while routine patterns warrant attenuation. We use a learnable gating mechanism to adjust residual contributions according to their predictive value for intention inference. Specifically, we derive gating coefficients for the orthogonally filtered residuals $\Delta^{\perp,\mathrm{in}}_{T}$ and $\Delta^{\perp,\mathrm{out}}_{T}$ using a linear layer followed by sigmoid activation, and then integrate these gated residuals with the original features:
$h^{\mathrm{in}}=X^{\mathrm{in}}_{T}+g^{\mathrm{in}}_{T}\cdot\Delta^{\perp,\mathrm{in}}_{T},\quad h^{\mathrm{out}}=X^{\mathrm{out}}_{T}+g^{\mathrm{out}}_{T}\cdot\Delta^{\perp,\mathrm{out}}_{T},$ (8)
where $g^{\mathrm{in}}_{T}$ and $g^{\mathrm{out}}_{T}$ are learned gating coefficients that selectively modulate causal signal contributions to enhance prediction robustness.
Adaptive Intention Encoding.
Beyond frame-level causal cues, holistic intent understanding requires global semantic reasoning. We extract a coarse intent distribution from the exterior summary $h^{\mathrm{out}}$ through softmax classification over $M$ predefined intention categories ($M$ denotes the total number of classes):
$\boldsymbol{\xi}=\text{softmax}(W_{\mathrm{int}}h^{\mathrm{out}})\in\mathbb{R}^{M}$ (9)
where $W_{\mathrm{int}}\in\mathbb{R}^{M\times D}$ linearly projects the $D$-dimensional exterior summary onto $M$ intention logits. This intention distribution is then re-embedded as a dense intention token $z_{\mathrm{intent}}=W_{\mathrm{proj}}\boldsymbol{\xi}\in\mathbb{R}^{D}$ that encodes the driving intention in continuous representation. The intention token serves as a global semantic anchor, providing top-down guidance across processing streams for consistent interpretation of ambiguous scenarios.
Table 1: Comparison of state-of-the-art methods on the Brain4Cars dataset using camera and additional sensor modalities (GPS, Map, Speed). The best results are highlighted in bold.
Method Camera GPS Map Speed Pr Re F1-score
IOHMM (Jain et al. 2015) 74.2 71.2 72.7
SDAE (Rekabdar and Mousas 2018) 71.9 74.8 73.3
AIO-HMM (Jain et al. 2015) 77.4 71.2 74.2
Deep CNN (Rekabdar and Mousas 2018) 78.0 77.5 77.7
FRNN-UL (Jain et al. 2016b) 82.2 75.9 78.9
FRNN-EL (Jain et al. 2016b) 84.5 77.1 80.6
FRNN-EL w/ 3D head pose (Jain et al. 2016b) 90.5 87.4 88.9
LSTM-GRU (Tonutti et al. 2019) 92.3 90.8 91.3
DCNN (Rekabdar and Mousas 2018) 91.8 92.5 92.1
CF-LSTM (Zhou et al. 2021) 92.0 92.3 92.1
Predictive-Bi-LSTM-CRF (Zhou et al. 2021) 92.4 94.7 93.6
Central (Zhu et al. 2024) 94.4 94.3 94.2
FedPRM (Zhu et al. 2024) 99.0 92.0 95.2
Gebert (Gebert et al. 2019) - - 81.7
Rong (Rong, Akata, and Kasneci 2020) - - 84.3
CEMFormer (Ma et al. 2023) - - 87.1
TIFN (Guo et al. 2023) 89.3 86.4 87.9
IDIPN (Liu et al. 2025) 94.2 94.9 94.5
CaSTFormer (Ours) 96.7 98.5 97.6
98.7 98.5 98.6
Feature Synthesis Network (FSN)
The CPE module provides a set of disentangled feature vectors corresponding to internal and external cues and a preliminary intent token. We further introduce the Feature Synthesis Network (FSN), which performs adaptive fusion of these features to construct a superior synthesized representation for predicting driving intention. By selectively emphasizing the most relevant information, the FSN module enhances the robustness of driving intention prediction. Each visual branch undergoes a residual nonlinear transformation via a dual-stage feedforward network with intermediate activation, which, combined with the speed feature $s$, yields the fused representations for the internal, external, and interaction streams:
$r_{\mathrm{in}}$ $=f_{\mathrm{in}}\bigl{(}[h_{\mathrm{in}},\,z_{\mathrm{intent}}]\bigr{)}+h_{\mathrm{in}},$ (10)
$r_{\mathrm{out}}$ $=f_{\mathrm{out}}\bigl{(}[h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{out}},$
$r_{\mathrm{ctx}}$ $=f_{\mathrm{ctx}}\bigl{(}[h_{\mathrm{in}},\,h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{in}}+h_{\mathrm{out}}$
where each $f_{\bullet}$ denotes a dual-stage feedforward mapping comprising two fully connected layers separated by a ReLU activation (FC–ReLU–FC). Let $\mathcal{C}=\{\mathrm{in},\mathrm{out},\mathrm{ctx}\}$. Each refined feature $r_{i}$ ($i\in\mathcal{C}$) is mapped to class logits $\ell_{i}$ and a corresponding confidence weight $w_{i}$, which adaptively controls each branch’s contribution:
$w_{i}=\frac{\exp(u_{i}^{\top}r_{i})}{\sum_{j\in\mathcal{C}}\exp(u_{j}^{\top}r_{j})},\quad\ell_{\mathrm{joint}}=\sum_{i\in\mathcal{C}}w_{i}\,(W_{i}\,r_{i}).$ (11)
Model Training
To address class imbalance and enhance sensitivity to rare intentions while promoting early prediction, we design a unified loss function that combines the average cross-entropy (CE) across complementary streams with an intention-prediction term:
$\mathcal{L}=\underbrace{\frac{1}{4}\sum_{i\in\ \mathcal{C}}\mathrm{CE}\bigl{(}\ell_{i},\,y\bigr{)}}_{\text{main loss}}\;+\;\underbrace{\alpha\,\mathrm{CE}\bigl{(}\ell_{\mathrm{intent}},\,y\bigr{)}}_{\text{intention loss}}$ (12)
where $\alpha$ is a tunable weight that governs the contribution of the intention‐prediction term. This unified objective integrates class-imbalance mitigation, multi-view fusion, and intention supervision within a cohesive framework.
Experiments
Data Preparation
Brain4Cars: The Brain4Cars dataset (Jain et al. 2016a) comprises exterior (480 $\times$ 720) and interior (1088 $\times$ 1920) videos of up to 5-second segments, refined to 594 valid events after excluding incomplete or unsynchronized samples. Each video is uniformly sampled to 150 frames, extracting the 5-second segment preceding the maneuver. Interior frames are cropped to 900 $\times$ 800, resized to 112 $\times$ 112, and encoded by a MobileFaceNet yielding 64-D features. Exterior frames undergo RAFT optical flow computation, are resized to 144 $\times$ 96, and then processed by ResNet-18 to produce 32-D features. Appending a smoothed speed signal yields 65-D (internal) and 33-D (external) vectors. These vectors are linearly projected, positionally encoded, and passed through a Transformer encoder to obtain temporal representations for CaSTFormer. The dataset spans highway and urban settings with five maneuver classes: straight, left turn, right turn, left lane change, and right lane change.
Implementation Details
Our proposed CaSTFormer was implemented by PyTorch and experiments were performed on a server with six NVIDIA RTX 2080 Ti GPUs. The model was trained end‐to‐end on Brain4Cars using the Adam optimizer (initial learning rate $1\times 10^{-3}$) for 160 epochs with a batch size of 16. During training, each input comprised a chunk of frames randomly sampled from the 5-second pre-maneuver segment; at inference, chunks were obtained via uniform sampling. The weight $\alpha$ in the unified loss was empirically set to 0.1. Model performance was evaluated using 5-fold cross-validation.
Evaluation Protocols
In driving intention prediction, straight driving is considered background, and only turns and lane changes are treated as target events. To evaluate our CaSTFormer model, we define the following metrics based on the predictions: true positives (TP: correctly predicted maneuvers), false positives (FP: maneuvers predicted incorrectly as another maneuver), false positive predictions (FPP: predicting a maneuver when none occurred), and missing predictions (MP: failing to detect an actual maneuver). Given the set of all behaviors $\mathcal{G}$ and target maneuvers $\mathcal{G}^{\prime}=\mathcal{G}\setminus\{\mathrm{straight}\}$, Precision (Pr), Recall (Re), and F1-score are computed as follows:
$\mathrm{Pr}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+FP_{m}+FPP_{m}},$ (13)
$\mathrm{Re}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+MP_{m}},\quad F_{1}=\frac{2*\mathrm{Pr}*\mathrm{Re}}{\mathrm{Pr}+\mathrm{Re}}.$
Comparison with State-of-the-art Methods
Table 2: Comparison of our CaSTFormer against other end-to-end methods on the Brain4Cars dataset, using internal and external streams, with F1-score (%) and parameters (M).
Method Inside Outside F1-score (%) Param. (M)
Gebert (Gebert et al. 2019) $\checkmark$ 81.7 85.3+162
$\checkmark$ 43.4 85.3+162
$\checkmark$ $\checkmark$ 73.2 170.5+162
Rong (Rong, Akata, and Kasneci 2020) $\checkmark$ 75.5 46.2+162
$\checkmark$ 66.4 5.4+162
$\checkmark$ $\checkmark$ 84.3 57.9+162
TIFN (Guo et al. 2023) $\checkmark$ $\checkmark$ 87.9 12.3+5.3
IDIPN (Liu et al. 2025) $\checkmark$ $\checkmark$ 94.5 11.75+5.3
CaSTFormer (Ours) $\checkmark$ $\checkmark$ 98.6 14.53+5.3
Table 1 provides a systematic comparison of both single- and multi-modal methods on the Brain4Cars dataset. Notably, our camera-only CaSTFormer variant achieves an F1-score of 97.6% (precision 96.7%, recall 98.5%), markedly surpassing all previous single-modality methods such as DCNN (92.1%) and CF-LSTM (92.1%). When enriched with speed information, CaSTFormer attains a new state-of-the-art F1-score of 98.6% (precision 98.7%, recall 98.5%), outperforming the best prior multi-modal model, FedPRM (95.2% F1), by 3.4%. These results demonstrate that CaSTFormer not only establishes a new performance standard but does so with fewer sensor inputs, highlighting its efficiency and robustness in driving intention prediction. Figure 4 presents the confusion matrices of our CaSTFormer and TIFN (Guo et al. 2023). CaSTFormer yields a sharper diagonal and substantially fewer off-diagonal entries, demonstrating its superior discrimination of similar maneuvers and reduced false predictions.
Table 3: F1-scores on the Brain4Cars dataset for evaluation on video segments truncated 1–4 s prior to action onset.
Method F1-score (%)
[-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
Rong (Rong, Akata, and Kasneci 2020) (in) 75.7 73.1 68.6 58.5 48.2
Rong (Rong, Akata, and Kasneci 2020) (out) 66.4 62.4 47.0 38.8 38.9
Rong (Rong, Akata, and Kasneci 2020) (both) 84.3 78.9 70.6 60.3 53.4
TIFN (Guo et al. 2023) 87.9 80.9 71.0 55.0 44.6
IDIPN (Liu et al. 2025) 94.5 84.1 74.2 62.0 55.4
CaSTFormer (Ours) 98.6 97.4 90.1 78.4 63.7
Detailed comparative results between our method and other end-to-end approaches on full-video inputs appear in Table 2. As the optical-flow algorithm lies outside the core prediction pipeline, its parameters are listed separately. Our model achieves superior recognition performance with only a marginal increase in model size, demonstrating its compact efficiency.
In addition to evaluating F1-scores on complete videos (–5 s to 0 s, where 0 s marks driver action), we assessed early-warning capability by truncating observation windows at –1 s, –2 s, –3 s, and –4 s. As shown in Table 3, prediction accuracy declines nearly linearly with shorter observations, highlighting increased uncertainty at longer forecast horizons. This result reflects the intrinsic trade-off between early intervention and predictive accuracy. Our CaSTFormer consistently achieves superior performance across all truncated settings, demonstrating its robustness in driving intention prediction.
Figure 4: The confusion matrix tested on Brain4cars dataset. Left is ours, right is the result of TIFN (Guo et al. 2023). The color deepens as the value increases.
Ablation Study
Effect of components in CaSTFormer.
To evaluate the contribution of each component in CaSTFormer, we conducted systematic ablation studies with results presented in Table 4. Using a dual-stream Transformer baseline (T-Base) that independently encodes internal and external features through simple concatenation, we progressively introduced the RSF, CPE and FSN modules to assess their individual impact on F1-score performance. We further explored various combined configurations to understand module interactions and their cumulative effects. The experiments reveal consistent performance improvements across all configurations, validating that each module makes meaningful contributions to driving intention prediction while their synergistic combination yields additive enhancement benefits.
Table 4: F1-scores on the Brain4Cars dataset between the dual-stream Transformer baseline and progressively augmented variants that introduce RSF, CPE and FSN.
Model F1‐score (%)
[-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
T-Base 95.8 94.2 85.4 73.7 63.2
T-Base+RSF 97.1 95.6 87.1 75.2 61.9
T-Base+CPE 97.0 95.4 86.9 74.9 61.1
T-Base+FSN 96.6 94.9 86.3 73.9 62.6
T-Base+RSF+CPE 97.4 95.7 87.4 75.9 60.3
T-Base+RSF+FSN 98.0 96.7 88.5 76.8 65.6
T-Base+CPE+FSN 97.8 96.4 88.0 76.2 62.4
T-Base+RSF+CPE +FSN(CaSTFormer) 98.6 97.4 90.1 78.4 63.7
Figure 5: F1‐scores (%) of different attention designs on the Brain4Cars dataset.
Effect of attention design in CaSTFormer.
To evaluate the effect of different attention mechanisms on driving intention prediction, we conducted ablation experiments on the Brain4Cars dataset comparing three attention schemes, as summarized in Table 5. The results indicate that our proposed Bidirectional Dependency Attention (BDA) more effectively captures spatio-temporal correlations between interior and exterior streams while suppressing noise, thereby demonstrating its superiority in isolating dynamic cues and enhancing overall model robustness.
Table 5: F1‐scores (%) of different the sequential correspondence orders in CaSTFormer on the Brain4Cars dataset.
Module Sequence F1‐score (%)
[-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
[…]
Effect of module sequencing in CaSTFormer.
[…]
Figure 6: F1‐scores (%) for different values of $\alpha$ in the loss function on the Brain4Cars dataset.
Effect of intention loss.
To assess the influence of the intention‑loss weight $\alpha$ in our composite objective, we conducted an ablation study by varying $\alpha$ to identify the optimal balance between core spatio‑temporal feature learning and dedicated intention supervision. The results in Fig. 6 show that setting $\alpha=0.10$ produces the highest F1‑scores across most anticipation horizons, because this value injects enough gradient from the intention loss to enrich the discriminative power of spatio‑temporal embeddings with causal cues while preserving the stability and generalization capacity of the primary classification objective. Smaller values of $\alpha$ under‑utilize intention supervision and fail to capture these pre‑action signals, whereas larger values allow the intention loss to dominate optimization, creating gradient conflicts that degrade both spatio‑temporal coherence and overall anticipation accuracy.
Conclusion
In this paper, we propose CaSTFormer, an interpretable prototype-driven causal spatio-temporal transformer for driving intention prediction. Our approach extracts dual‐stream interior and exterior features and processes them through a structured pipeline to align and fuse multi‐modal representations. Extensive experiments on the Brain4Cars dataset demonstrate that CaSTFormer achieves state‐of‐the‐art prediction accuracy with only a marginal increase in model size, validating its efficiency and suitability for real‐time driver assistance.

### v2

@@
-CaSTFormer: Causal Spatio‐Temporal Transformer for Driving Intention Prediction
+CaTFormer: Causal Temporal Transformer with Dynamic Contextual Fusion for Driving Intention Prediction
 Abstract
@@
 It serves as a cornerstone for achieving high-level autonomous driving.
-However, current approaches remain inadequate for accurately modeling the complex spatio-temporal interdependencies and the unpredictable variability of human driving behavior.
-To address these challenges, we propose CaSTFormer, a Causal Spatio-Temporal Transformer to explicitly model causal interactions between driver behavior and environmental context for robust intention prediction.
-Specifically, CaSTFormer introduces a novel Reciprocal Shift Fusion (RSF) mechanism for precise temporal alignment of internal and external feature streams, a Causal Pattern Extraction (CPE) module that systematically eliminates spurious correlations to reveal authentic causal dependencies, and an innovative Feature Synthesis Network (FSN) that adaptively synthesizes these purified representations into coherent spatio-temporal inferences.
-We evaluate the proposed CaSTFormer on the public Brain4Cars dataset, and it achieves state-of-the-art performance.
-It effectively captures complex causal spatio-temporal dependencies and enhances both the accuracy and transparency of driving intention prediction.
+However, current approaches remain inadequate for accurately modeling the complex spatiotemporal interdependencies and the unpredictable variability of human driving behavior.
+To address these challenges, we propose CaTFormer, a causal Temporal Transformer that explicitly models causal interactions between driver behavior and environmental context for robust intention prediction.
+Specifically, CaTFormer introduces a novel Reciprocal Delayed Fusion (RDF) mechanism for precise temporal alignment of interior and exterior feature streams, a Counterfactual Residual Encoding (CRE) module that systematically eliminates spurious correlations to reveal authentic causal dependencies, and an innovative Feature Synthesis Network (FSN) that adaptively synthesizes these purified representations into coherent temporal representations.
+Experimental results demonstrate that CaTFormer attains state-of-the-art performance on the Brain4Cars dataset.
+It effectively captures complex causal temporal dependencies and enhances both the accuracy and transparency of driving intention prediction.
+Codehttps://github.com/srwang0506/CaTFormer
 Introduction
-Driver intention prediction is critical in autonomous driving systems, playing a significant role in effectively mitigating risks and enhancing driving safety.
-By forecasting potential outcomes several seconds in advance, the system can alert the driver proactively or initiate evasive maneuvers, further increasing its safety capabilities.
-Typically, the intention prediction relied on the extraction and fusion of visual features to perform simple intention prediction (Huang et al.
+Driver intention prediction is crucial for autonomous driving systems, as it effectively mitigates risks and enhances driving safety.
+By forecasting potential outcomes several seconds in advance, the system can proactively alert the driver or initiate evasive maneuvers, significantly improving its safety capabilities.
+Initially, intention prediction primarily relied on the extraction and fusion of visual features for basic predictions (Huang et al.
 2022).
-The development of sensor technology has enabled the use of multi-modal information, including GPS coordinates, vehicle speed, map data, and driver head pose (Hu et al.
+However, advancements in sensor technology have enabled the incorporation of multi-modal information, such as GPS coordinates, vehicle speed, map data, and driver head pose (Hu et al.
 2021; Mo et al.
-2023; Li, Zhao, and Wang 2022; Wu et al.
+2023; Li et al.
+2022; Wu et al.
 2023).
-By leveraging complex models for feature extraction, fusion, and prediction, the performance of driver intention prediction has been significantly improved (Sui et al.
+By leveraging complex models for feature extraction, fusion, and prediction, the performance of driver intention prediction has significantly improved (Sui et al.
 2021; Guo et al.
-2023; Liu, Wu, and Wang 2023; Gao et al.
+2023; Liu et al.
+2023; Gao et al.
 2023).
 Figure 1: Comparison between previous driving intention prediction methods and ours.
-(a) is an LSTM–GRU–based framework in which internal and external streams are processed independently and then concatenated for final fusion.
-(b) is our CaSTFormer, a Transformer–based model that performs dynamic causal fusion of the dual streams while integrating intention priors.
-By jointly capturing global and local dependencies and leveraging intrinsic cross-stream interactions, our method surpasses existing approaches.
-Despite the progress made in driver intention prediction, the rich multi-modal data has not been fully exploited.
-Most existing work merely concatenates or linearly aggregates this information, as illustrated in Fig.
+(a) is an LSTM-GRU framework processing interior and exterior streams independently before concatenation.
+(b) is our CaTFormer, a Transformer-based model enabling dynamic causal fusion of dual streams with integrated intention priors.
+Through joint modeling of global-local dependencies and cross-stream interactions, our approach outperforms existing methods.
+Despite advancements in driver intention prediction, the rich multi-modal data remains underutilized.
+Most existing studies simply concatenate or linearly aggregate this information, as shown in Fig.
 1 (a).
-However, as the controller of the vehicle, changes in the driver’s state subsequently influence the vehicle’s driving status, revealing a tight dependency.
-Therefore, we propose to explicitly model the causal relationship between the driver and the environment, emphasizing the decisive impact of this causality on the prediction task, as shown in Fig 1 (b).
+However, given that the driver controls the vehicle, changes in their state directly influence the vehicle’s driving status, indicating a strong dependency.
+Therefore, we propose to explicitly model the causal relationship between the driver and the environment, highlighting this causality’s decisive impact on the prediction task, as illustrated in Fig.
+1 (b).
 Specifically, we adopt a Transformer-based architecture and introduce three sequential components to enhance the causal modeling, encode the driver’s intention, and fuse the multi-dimensional features.
-First, we introduce a Reciprocal Shift Fusion (RSF) module that cross-fuses internal and external features through a shifting mechanism, explicitly establishing a temporal dependency between the two feature streams.
-As the fused features may contain considerable casual noise, we further devise a Causal Pattern Extraction (CPE) module to filter out such noise to obtain a clearer causal representation.
-Finally, we utilize a Feature Synthesis Network (FSN) that employs a gating mechanism to integrate the in-cabin, out-of-cabin, and interaction representations, enabling the modeling of both local and global causal structures.
+First, we introduce a Reciprocal Delayed Fusion (RDF) module that cross-fuses interior and exterior features through a shifting mechanism, explicitly establishing a temporal dependency between the two feature streams.
+As the fused features may contain considerable causal noise, we further devise a Counterfactual Residual Encoding (CRE) module to filter out such noise to obtain a more explicit causal representation.
+Finally, we utilize a Feature Synthesis Network (FSN) that employs a gating mechanism to integrate the interior, exterior, and interaction representations, enabling the modeling of both local and global causal structures.
 Our main contributions are as follows:
-  • We propose CaSTFormer, an efficient Transformer-based framework for driving-intent prediction that integrates temporal dependency modeling, causal pattern enrichment, and adaptive multi-view fusion within a unified end-to-end architecture.
-  • CaSTFormer disentangles spatio‐temporal interactions of interior and exterior streams, subsequently isolates genuine causal contributions through counterfactual attention subtraction and adaptively fuses heterogeneous features, thereby mitigating spurious correlations and improving robustness under complex driving conditions.
-  • Extensive evaluation on the Brain4Cars dataset demonstrates that CaSTFormer demonstrates superior performance in driving intention prediction in both highway and urban scenarios.
+  • We propose CaTFormer, an efficient Transformer-based framework for driving intention prediction that embeds causal spatio‑temporal reasoning with adaptive multi‑view fusion in a unified end‑to‑end architecture.
+  • Through dual-stream reciprocal delayed fusion, CaTFormer explicitly captures dependencies across interior and exterior streams, isolates genuine causal effects via counterfactual attention subtraction, and adaptively integrates complementary visual cues, effectively enhancing robustness under complex driving conditions.
+  • Extensive evaluation on the Brain4Cars dataset demonstrates that CaTFormer demonstrates superior performance in driving intention prediction in both highway and urban scenarios.
 Related Work
-In the early stages of research on driving intention prediction, some studies focused on learning spatio-temporal features directly from raw video data.
-For instance, a combination of 3D Convolutional Neural Networks and Long Short-Term Memory networks was utilized to predict driving maneuvers directly from video streams (Gebert et al.
+In the early stages of driving intention prediction research, studies primarily focused on learning spatiotemporal representations directly from raw video, often employing 3D CNN-LSTM architectures for maneuver prediction (Gebert et al.
 2019).
-To achieve a more comprehensive basis for decision-making, multi-modal data was introduced to improve the accuracy of intention prediction.
-In one such study, the in-cabin driver’s view was combined with the external traffic scene, using a Convolutional Long Short-Term Memory network to fuse and analyze features from both views (Rong, Akata, and Kasneci 2020).
-However, their contextual capacity is limited, making it difficult to process long-range, non-consecutive events.
+While some later work attempted to enrich context by fusing interior and exterior streams via convolutional LSTMs for a more comprehensive decision-making basis (Rong et al.
+2020), both approaches struggled with limited capacity for modeling long-range, non-consecutive dependencies.
 Inspired by human cognitive processes, TIFN (Guo et al.
-2023) proposed a State Update Unit to incorporate the influence of environmental information into driver state modeling, and extracted semantic segmentation features to provide clear cues affecting driver attention.
-Meanwhile, another approach treated intention prediction as a sequence labeling task, further capturing the contextual dependencies of driving behavior by combining a Bidirectional Long Short-Term Memory network with a Conditional Random Field (Zhou et al.
+2023) introduced a state update unit (STU) to integrate environmental context into driver state modeling and extract semantic segmentation features as attention cues.
+Similarly, another study framed intention prediction as a sequence-labeling task, combining bidirectional LSTMs with a conditional random field to capture the contextual dependencies of driving behaviors (Zhou et al.
 2021).
-Although these models fuse multi-source information, they typically learn the correlations between features implicitly, without explicitly decoupling and reasoning about the mutual influences between different information modalities.
-Furthermore, personalized differences among drivers were also taken into consideration as a significant source of model error.
-Early work used a Domain-Adversarial Recurrent Neural Network (DA-RNN) to learn driver-agnostic, domain-invariant features to enhance the model’s generalization capabilities (Tonutti et al.
-2019).
-Recently, IDIPN (Liu et al.
-2025) utilized Inverse Reinforcement Learning to extract unique driving preferences from each driver’s historical data, thereby correcting the predictions of a general-purpose model, while FedPRM (Zhu et al.
-2024) achieved this by using a federated learning framework to directly construct personalized models for each driver.
-The Transformer architecture, with its unique attention mechanism, excels at capturing long-range dependencies and global interaction information in temporal data (Liu et al.
+Although these methods incorporate multi-source information, they typically learn inter-modal feature correlations implicitly, lacking explicit disentanglement and reasoning of their interactions.
+Moreover, to enhance model generalization, existing studies have developed personalized prediction models using techniques like domain-adversarial RNN (Tonutti et al.
+2019), inverse reinforcement learning (Liu et al.
+2025), and a federated learning framework (Zhu et al.
 2024).
-CemFormer (Ma et al.
-2023) integrates data from both in-cabin and external cameras, learning a unified cross-view representation via a spatio-temporal tansformer to infer driver intention directly from their behavior.
-However, such methods primarily rely on the inherent structure of the transformer to interpret the dependencies.
-Meanwhile, a non-autoregressive transformer with hybrid attention has been utilized to simultaneously capture the temporal dynamics of a single vehicle, as well as the interactions among multiple vehicles (Jiang et al.
+The Transformer architecture excels at capturing long-range dependencies and global interaction information in temporal data through its unique attention mechanism (Liu et al.
+2024).
+Building on this, CemFormer (Ma et al.
+2023) integrates data from both interior and exterior cameras, learning a unified cross-view representation via a spatio-temporal Transformer to infer driver intention directly from their behavior.
+However, these methods primarily rely on the Transformer’s inherent structure to interpret dependencies.
+Meanwhile, a non-autoregressive Transformer with hybrid attention has been employed to simultaneously capture the temporal dynamics of a single vehicle and interactions among multiple vehicles (Jiang et al.
 2024).
 DriveTransformer (Jia et al.
-2025) further constructs a unified end-to-end framework to handle perception, prediction, and planning tasks in parallel.
-This treats intention prediction as an integrated component of the overall scalable system, but may dilute the model’s focus on driver-intention features.
-Besides, some early works have attempted to introduce causal inference methods into the intention prediction task.
-One study constructed a causal model to learn the temporal relationships of invariant representations from driving data, thereby achieving domain generalization (Hu et al.
+2025) further establishes a unified end-to-end framework to handle perception, prediction, and planning tasks in parallel.
+While this approach treats intention prediction as an integrated component of a scalable system, it may dilute the model’s focus on specific driver-intention features.
+Besides, some early works also integrated causal inference methods into intention prediction.
+For instance, one study built a causal model to capture temporal relationships within invariant representations from driving data, aiming for domain generalization (Hu et al.
 2022).
-Another adopted a driver-centric approach, formulating the task of risk object identification as a causal problem and proposing a two-stage framework based on causal inference (Li, Chan, and Chen 2020).
-However, the scope of scenarios and conditions covered by these methods is relatively limited.
-Apart from all the prior work, our method employs a dual-stream transformer.
-Building upon an encoding of driver intention, it explicitly establishes both local and global causal dependencies between in-cabin and external factors.
-This approach, through the fusion of multi-dimensional features, enables precise intention prediction to address increasingly complex driving scenarios.
-Figure 2: Overview of the CaSTFormer pipeline.
-After data preprocessing, external optical flow is encoded by ResNet-18 and interior images by MobileFaceNet to produce dual-stream feature sequences.
-These are then fed into three core modules: (a) Reciprocal Shift Fusion (RSF) for temporal feature integration; (b) Causal Pattern Extraction (CPE) for causal enhancement and intent embedding; and (c) Feature Synthesis Network (FSN) for dynamic fusion of complementary internal, external and interaction views to yield the final driving intention prediction.
+Another adopted a driver-centric approach, framing risk object identification as a causal inference problem and introducing a two-stage causal framework (Li et al.
+2020).
+However, these methods typically cover a relatively limited scope of scenarios and conditions.
+Building on prior work, we propose a dual‑stream Transformer architecture that leverages a learned intention embedding to explicitly capture both local and global causal dependencies between in‑cabin and external modalities.
+Through systematic fusion of these multi‑dimensional feature representations, the model achieves highly accurate intention prediction in complex driving environments.
+Figure 2: Overview of the CaTFormer pipeline.
+After data preprocessing, exterior optical flow is encoded by ResNet-18 and interior images by MobileFaceNet to produce dual-stream feature sequences.
+These are then fed into three core modules: (1) Reciprocal Delayed Fusion (RDF) for temporal feature integration; (2) Counterfactual Residual Encoding (CRE) for causal enhancement and intention embedding; and (3) Feature Synthesis Network (FSN) for dynamic fusion of complementary interior, exterior and interaction views to yield the final driving intention prediction.
 Method
 As illustrated in Fig.
-2, CaSTFormer takes as input a bi-stream image sequence =$\mathcal{I}=\{(I^{\mathrm{out}}_{b,t},I^{\mathrm{in}}_{b,t})\}_{b,t=1}^{B,T}$ comprising $B$ samples, each consisting of $T$ frames from two synchronized streams (outside and inside views), where $b$ and $t$ index the sample and temporal frame, respectively.
-Feature extractors separately encode each stream into features $F^{\mathrm{out}},F^{\mathrm{in}}\in\mathbb{R}^{B\times T\times D}$ ($D$ denotes the dimensionality of each encoded feature vector).
-The obtained features pass through a Reciprocal Shift Fusion (RSF) module for temporal causality modeling via our proposed bi-stream attention, followed by a Causal Pattern Extraction (CPE) module to inject learnable causal representations.
+2, CaTFormer processes a bi‐stream image sequence $\mathcal{I}=\bigl\{(I^{\mathrm{out}}_{b,t},I^{\mathrm{in}}_{b,t})\bigr\}_{b=1,t=1}^{B,T}$ of $B$ synchronized frame pairs over $T$ time steps, where $b$ and $t$ index the sample and temporal frame, respectively.
+Feature extractors encode each stream into features $F^{\mathrm{out}},F^{\mathrm{in}}\in\mathbb{R}^{B\times T\times D}$, where $D$ denotes the dimensionality of each encoded feature vector.
+The features pass through a Reciprocal Delayed Fusion (RDF) module for temporal causality modeling via our proposed bi-stream attention, followed by a Counterfactual Residual Encoding (CRE) module to inject learnable causal representations.
 Finally, visual features ($z_{\mathrm{in}},z_{\mathrm{out}}$ and interaction features $z_{\mathrm{ctx}}$) are adaptively fused by a Feature Synthesis Network (FSN) into a joint prediction $\ell_{\mathrm{joint}}$.
-Reciprocal Shift Fusion (RSF)
-On multi‐lane highways, cameras concurrently record the external traffic scene and the driver’s internal state.
-Recognizing the mutual influences between environmental context and driver behavior, our dual‐stream architecture explicitly captures their bidirectional interactions by jointly modeling both feature streams.
-To enforce exterior-first precedence, we introduce a temporal delay mechanism in the Key and Value sequences.
-Specifically, at time step $t$, the attention mechanism accesses only information from the preceding frame $t-1$.
+Reciprocal Delayed Fusion (RDF)
+On multi‐lane highways, cameras concurrently record the exterior traffic scene and the driver’s interior state.
+Motivated by bidirectional dependencies between environmental context and driving behavior, our dual‐stream architecture explicitly models their interactions by jointly processing both feature streams.
+To model inter-frame temporal precedence, we introduce a temporal delay mechanism in the Key and Value sequences.
+Specifically, at time step $t$, the attention mechanism accesses only information from the preceding frame t1t-1.
 Concretely, we define the delayed feature
-[…]
+$\hat{F}_{b,t}=F_{b,t-1}\,\mathbf{1}_{\{t>1\}},\quad\mathbf{1}_{\{t>1\}}=\begin{cases}1,&t>1,\\ 0,&t=1.\end{cases}$ (1)
 applied separately to $F^{\mathrm{out}}$ and $F^{\mathrm{in}}$.
 Bidirectional Dependency Attention (BDA).
-Under a strict single-frame delay constraint, BDA enriches each frame’s representation by fusing internal and external contexts from the immediately preceding timestep.
-The current internal and external features attend bidirectionally to their one-frame–delayed counterparts, capturing both temporal coherence and cross-stream coupling.
-To model diverse associations efficiently, we project into $H$ parallel attention heads and aggregate their outputs through concatenation and a final linear mapping:
+Under a strict single-frame delay constraint, BDA enriches each frame’s representation by fusing interior and exterior contexts from the immediately preceding timestep.
+The current interior and exterior features attend bidirectionally to their one-frame delayed counterparts, capturing both temporal coherence and cross-stream coupling.
+To model diverse associations efficiently, we project into $H$ parallel attention heads (in our experiments, $H=8$) and aggregate their outputs through concatenation and a final linear mapping:
 $\mathrm{BDA}(Q,K,V)=\Bigl[\mathrm{softmax}\bigl(\tfrac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}\bigr)V_{i}\Bigr]_{i=1}^{H}W^{O},$ (2)
-where $[\cdot]_{i=1}^{H}$ denotes head-wise concatenation and $W^{O}$ restores the original feature dimension.
-Figure 3 illustrates the detailed bidirectional query–key–value fusion, showing how the internal and external streams are jointly refined.
-Figure 3: Illustration of Bidirectional Dependency Attention (BDA) where buffered internal and external features are mutually cross-attended to produce dynamically enhanced current-frame representations.
+where $d_{k}$ is the key dimension, $\cdot]_{i=1}^{H}$ denotes concatenation across heads, and $W^{O}$ restores the original feature size.
+Fig.
+3 shows the bidirectional query-key-value fusion, highlighting how interior and exterior streams are jointly updated.
+Figure 3: Illustration of Bidirectional Dependency Attention (BDA), where buffered interior and exterior features cross-attend to enhance current-frame representations.
 Channel-wise gating.
-Although BDA produces a richly fused representation $H_{b,t}$ at each spatial–temporal location, individual channels may still carry irrelevant or noisy information.
-Hence, we employ two successive channel‐wise gating layers that adaptively emphasize informative channels and suppress spurious responses:
-$g_{b,t}$ $=\sigma\bigl{(}W_{2}\,(\mathrm{ReLU}(W_{1}H_{b,t}+b_{1}))+b_{2}\bigr{)}$ (3)
-$\widetilde{H}_{b,t}$ $=g_{b,t}\,\odot\,H_{b,t}$
-where $\odot$ denotes the Hadamard product between vectors.
+Although BDA generates a fused representation $H_{b,t}$ at each spatial-temporal point, some channels may still carry noise or irrelevant signals.
+Thus, we apply two channel-wise gating layers to adaptively enhance informative features and suppress spurious ones:
+$g_{b,t}$ $=\sigma\bigl(W_{2}\,(\mathrm{ReLU}(W_{1}H_{b,t}+b_{1}))+b_{2}\bigr),$ (3)
+$\widetilde{H}_{b,t}$ $=g_{b,t}\,\odot\,H_{b,t},$
+where $\sigma(\cdot)$ denotes the sigmoid activation function and $\odot$ denotes the Hadamard product between vectors.
 Normalization and regularization.
 For numerical stability and to guard against overfitting, the following operations are applied to the gated outputs $\widetilde{H}_{b,t}$:
-$R(x)=x\;\odot\;\frac{s}{\sqrt{\frac{1}{D}\sum_{d=1}^{D}x_{d}^{2}+\epsilon}}$ (4)
-${X}_{b,t}=\mathrm{Dropout}\bigl{(}R(\widetilde{H}_{b,t})\bigr{)}$ (5)
-where $R(\cdot)$ denotes the Root-mean-square normalization function, $s$ is a learnable scaling parameter, $\epsilon$ is a small constant to ensure numerical stability, and $D$ is the dimensionality of the feature vector $x$, yielding channel-calibrated representations across both internal and external feature streams.
-Causal Pattern Extraction (CPE)
-Conventional intention prediction architectures aggregate heterogeneous interior and exterior cues under an implicit correlation assumption and thus often mistake coincidental patterns for true decision drivers.
-To overcome this limitation, our CPE module contrasts observed and counterfactual cross‐stream attentions to disentangle direct causal contributions, then selectively amplifies only those residuals that genuinely influence driving intent, resulting in improved robustness and generalization in safety‐critical scenarios.
-Specifically, CPE takes the bidirectionally fused interior and exterior features ${X}_{b,T}^{\mathrm{in}},{X}_{b,T}^{\mathrm{out}}\in\mathbb{R}^{T\times B\times D}$ as inputs to perform this causal reasoning process.
+$R(x)=x\;\odot\;\frac{s}{\sqrt{\frac{1}{D}\sum_{d=1}^{D}x_{d}^{2}+\epsilon}},$ (4)
+${X}_{b,t}=\mathrm{Dropout}\bigl(R(\widetilde{H}_{b,t})\bigr),$ (5)
+where $R(\cdot)$ denotes the Root-mean-square normalization function, $s$ is a learnable scaling parameter, $\epsilon$ is a small constant to ensure numerical stability, and $D$ is the dimensionality of the feature vector $x$, yielding channel-calibrated representations across both interior and exterior feature streams.
+Counterfactual Residual Encoding (CRE)
+Conventional intention prediction architectures aggregate heterogeneous interior and exterior cues under an implicit correlation assumption and thus often mistake coincidental patterns for proper decision drivers.
+To overcome this limitation, our CRE module contrasts observed and counterfactual cross‐stream attentions to disentangle direct causal contributions.
+We then selectively amplify only those residuals that genuinely influence driving intention, resulting in improved robustness and generalization in safety‐critical scenarios.
+Specifically, CRE takes the bidirectionally fused interior and exterior features ${X}_{b,T}^{\mathrm{in}},{X}_{b,T}^{\mathrm{out}}\in\mathbb{R}^{T\times B\times D}$ as inputs to perform this causal reasoning process.
 Direct causal effect.
 At each time step, we compute two attention distributions.
-We first calculate observed attention $A^{\mathrm{obs}}_{\mathrm{in},t}$ using actual exterior features, and then generate counterfactual attention $A^{\mathrm{cf}}_{\mathrm{in},t}$ by replacing all exterior features with their temporal mean $\bar{X}^{\mathrm{out}}=\frac{1}{TB}\sum_{u=1}^{T}\sum_{b=1}^{B}X^{\mathrm{out}}_{u,b}$, which serves as a neutral baseline that removes environmental variations.
-The difference $\Delta^{\mathrm{in}}_{t}$ between these two distributions quantifies the direct causal influence of external context on interior representations.
+We first calculate observed dependency attention $A^{\mathrm{obs}}_{\mathrm{in},t}$ using actual exterior features, and then generate counterfactual dependency attention $A^{\mathrm{cf}}_{\mathrm{in},t}$ by replacing all exterior features with their temporal mean $\bar{X}^{\mathrm{out}}=\frac{1}{TB}\sum_{u=1}^{T}\sum_{b=1}^{B}X^{\mathrm{out}}_{u,b}$, which serves as a neutral baseline that removes environmental variations.
+The difference $\Delta^{\mathrm{in}}_{t}$ between these two distributions quantifies the direct causal attention of exterior context on interior representations.
 Specifically, for $t=1,\dots,T$, we obtain
-$A^{\mathrm{obs}}_{\mathrm{in},t}$ $=\mathcal{A}\bigl{(}X^{\mathrm{in}}_{t},\,X^{\mathrm{out}}_{\leq t-1},\,X^{\mathrm{out}}_{\leq t-1}\bigr{)},$ (6)
+$A^{\mathrm{obs}}_{\mathrm{in},t}$ $=\mathrm{BDA}\bigl(X^{\mathrm{in}}_{t},\,X^{\mathrm{out}}_{t-1},\,X^{\mathrm{out}}_{t-1}\bigr),$ (6)
 $A^{\mathrm{cf}}_{\mathrm{in},t}$ $=\mathcal{A}\bigl(X^{\mathrm{in}}_{t},\,\bar{X}^{\mathrm{out}},\,\bar{X}^{\mathrm{out}}\bigr),$
@@
 Similarly, $\Delta^{\mathrm{out}}_{t}$ is defined by interchanging the two streams.
-We further orthogonalize each causal residual against the global baseline vector $\bar{X}$ to ensure that the identified causal patterns reflect true intent-relevant dependencies rather than dataset-specific biases.
+We further orthogonalize each causal residual against the global baseline vector $\bar{X}$ to ensure that the identified causal patterns reflect true intention-relevant dependencies rather than dataset-specific biases.
 Formally,
-$\Delta_{t}^{\perp}=\Delta_{t}-\frac{\langle\Delta_{t},\,\bar{X}\rangle}{\|\bar{X}\|^{2}+\varepsilon}\,\cdot\bar{X},$ (7)
-where $\varepsilon$ is a small constant to ensure numerical stability.
-The orthogonal projection produces $\Delta^{\perp}_{t}$ that captures context-deviating residuals while removing baseline-aligned components.
-We retain the final-step residual $\Delta^{\perp}_{T}$ as the decision-relevant signal for downstream modules, ensuring that subsequent fusion operates exclusively on temporally salient causal information.
-Dynamic Causal Gating.
+$\Delta_{t}^{\perp}=\Delta_{t}-\frac{\Delta_{t}^{\top}\,\bar{X}}{\|\bar{X}\|^{2}+\varepsilon}\,\cdot\bar{X},$ (7)
+where $\varepsilon$ ensures numerical stability.
+The orthogonal projection yields $\Delta^{\perp}_{t}$ by removing baseline-aligned components.
+We retain $\Delta^{\perp}_{T}$ as the decision-relevant signal, guiding downstream fusion toward temporally salient causal cues.
+Dynamic residual gating.
 The causal relevance of residuals differs across driving scenarios, as critical maneuvers merit amplification, while routine patterns warrant attenuation.
@@
 Adaptive intention encoding.
-Beyond frame-level causal cues, holistic intent understanding requires global semantic reasoning.
-We extract a coarse intent distribution from the exterior summary $h^{\mathrm{out}}$ through softmax classification over $M$ predefined intention categories ($M$ denotes the total number of classes):
+Beyond frame-level causal cues, holistic intention understanding requires global semantic reasoning.
+We extract a coarse intention distribution from the exterior summary $h^{\mathrm{out}}$ through softmax classification over $M$ predefined intention categories ($M$ denotes the total number of classes):
 $\boldsymbol{\xi}=\text{softmax}(W_{\mathrm{int}}h^{\mathrm{out}})\in\mathbb{R}^{M}$ (9)
-where $W_{\mathrm{int}}\in\mathbb{R}^{M\times D}$ linearly projects the $D$-dimensional exterior summary onto $M$ intention logits.
-This intention distribution is then re-embedded as a dense intention token $z_{\mathrm{intent}}=W_{\mathrm{proj}}\boldsymbol{\xi}\in\mathbb{R}^{D}$ that encodes the driving intention in continuous representation.
+where $W_{\mathrm{int}}\in\mathbb{R}^{M\times D}$ linearly projects the $D$-dimensional exterior summary onto $M$ logits.
+This intention distribution is then re-embedded as an intention token $z_{\mathrm{intent}}\in\mathbb{R}^{D}$ that encodes the driving intention in a continuous representation.
 The intention token serves as a global semantic anchor, providing top-down guidance across processing streams for consistent interpretation of ambiguous scenarios.
-Table 1: Comparison of state-of-the-art methods on the Brain4Cars dataset using camera and additional sensor modalities (GPS, Map, Speed). The best results are highlighted in bold.
 Method Camera GPS Map Speed Pr Re F1-score
@@
 Gebert (Gebert et al. 2019) - - 81.7
-Rong (Rong, Akata, and Kasneci 2020) - - 84.3
+Rong (Rong et al. 2020) - - 84.3
 CEMFormer (Ma et al. 2023) - - 87.1
@@
 IDIPN (Liu et al. 2025) 94.2 94.9 94.5
-CaSTFormer (Ours) 96.7 98.5 97.6
+CaTFormer (Ours) 96.7 98.5 97.6
 98.7 98.5 98.6
+Table 1: Comparison of state-of-the-art methods on the Brain4Cars dataset using camera and additional sensor modalities (GPS, Map, Speed). The best results are highlighted in bold.
 Feature Synthesis Network (FSN)
-The CPE module provides a set of disentangled feature vectors corresponding to internal and external cues and a preliminary intent token.
+The CRE module provides a set of disentangled feature vectors corresponding to interior and exterior cues and a preliminary intention token.
 We further introduce the Feature Synthesis Network (FSN), which performs adaptive fusion of these features to construct a superior synthesized representation for predicting driving intention.
 By selectively emphasizing the most relevant information, the FSN module enhances the robustness of driving intention prediction.
-Each visual branch undergoes a residual nonlinear transformation via a dual-stage feedforward network with intermediate activation, which, combined with the speed feature $s$, yields the fused representations for the internal, external, and interaction streams:
-$r_{\mathrm{in}}$ $=f_{\mathrm{in}}\bigl{(}[h_{\mathrm{in}},\,z_{\mathrm{intent}}]\bigr{)}+h_{\mathrm{in}},$ (10)
-$r_{\mathrm{out}}$ $=f_{\mathrm{out}}\bigl{(}[h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{out}},$
-$r_{\mathrm{ctx}}$ $=f_{\mathrm{ctx}}\bigl{(}[h_{\mathrm{in}},\,h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{in}}+h_{\mathrm{out}}$
-where each $f_{\bullet}$ denotes a dual-stage feedforward mapping comprising two fully connected layers separated by a ReLU activation (FC–ReLU–FC).
+Each visual branch undergoes a residual nonlinear transformation via a dual-stage feedforward network with intermediate activation, which, combined with the speed feature $s$, yields the fused representations for the interior, exterior, and interaction streams:
+$r_{\mathrm{in}}$ $=f_{\mathrm{in}}\bigl([h^{\mathrm{in}},\,z_{\mathrm{intent}}]\bigr)+h^{\mathrm{in}},$ (10)
+$r_{\mathrm{out}}$ $=f_{\mathrm{out}}\bigl([h^{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr)+h^{\mathrm{out}},$
+$r_{\mathrm{ctx}}$ $=f_{\mathrm{ctx}}\bigl([h^{\mathrm{in}},\,h^{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr)+h^{\mathrm{in}}+h^{\mathrm{out}}$
+where each $f_{\cdot}$ denotes a dual-stage feedforward mapping comprising two fully connected layers separated by a ReLU activation (FC–ReLU–FC).
 Let $\mathcal{C}=\{\mathrm{in},\mathrm{out},\mathrm{ctx}\}$.
@@
 To address class imbalance and enhance sensitivity to rare intentions while promoting early prediction, we design a unified loss function that combines the average cross-entropy (CE) across complementary streams with an intention-prediction term:
-$\mathcal{L}=\underbrace{\frac{1}{4}\sum_{i\in\ \mathcal{C}}\mathrm{CE}\bigl{(}\ell_{i},\,y\bigr{)}}_{\text{main loss}}\;+\;\underbrace{\alpha\,\mathrm{CE}\bigl{(}\ell_{\mathrm{intent}},\,y\bigr{)}}_{\text{intention loss}}$ (12)
-where $\alpha$ is a tunable weight that governs the contribution of the intention‐prediction term.
+$\mathcal{L}=\underbrace{\frac{1}{4}\sum_{i\in\mathcal{H}}\mathrm{CE}\bigl(\ell_{i},\,y\bigr)}_{\text{main loss}}\;+\;\underbrace{\alpha\,\mathrm{CE}\bigl(\ell_{\mathrm{intent}},\,y\bigr)}_{\text{intention loss}}$ (12)
+where $\mathcal{H}=\{\mathrm{in},\mathrm{out},\mathrm{ctx},\mathrm{joint}\}$ denotes the four stream-level heads, and $\alpha$ controls the weight of the intention supervision term.
 This unified objective integrates class-imbalance mitigation, multi-view fusion, and intention supervision within a cohesive framework.
@@
 Interior frames are cropped to 900 $\times$ 800, resized to 112 $\times$ 112, and encoded by a MobileFaceNet yielding 64-D features.
-Exterior frames undergo RAFT optical flow computation, are resized to 144 $\times$ 96, and then processed by ResNet-18 to produce 32-D features.
-Appending a smoothed speed signal yields 65-D (internal) and 33-D (external) vectors.
-These vectors are linearly projected, positionally encoded, and passed through a Transformer encoder to obtain temporal representations for CaSTFormer.
+Exterior frames undergo Farneback optical flow computation, are resized to 144 $\times$ 96, and then processed by ResNet-18 to produce 32-D features.
+Appending a smoothed speed signal yields 65-D (interior) and 33-D (exterior) vectors.
+These vectors are linearly projected, positionally encoded, and passed through a Transformer encoder to obtain temporal representations for CaTFormer.
 The dataset spans highway and urban settings with five maneuver classes: straight, left turn, right turn, left lane change, and right lane change.
+Method In Out F1 (%) Param. (M)
+Gebert (Gebert et al. 2019) $\checkmark$ 81.7 85.3+162
+$\checkmark$ 43.4 85.3+162
+$\checkmark$ $\checkmark$ 73.2 170.5+162
+Rong (Rong et al. 2020) $\checkmark$ 75.5 46.2+162
+$\checkmark$ 66.4 5.4+162
+$\checkmark$ $\checkmark$ 84.3 57.9+162
+TIFN (Guo et al. 2023) $\checkmark$ $\checkmark$ 87.9 12.3+5.3
+IDIPN (Liu et al. 2025) $\checkmark$ $\checkmark$ 94.5 11.75+5.3
+CaTFormer (Ours) $\checkmark$ $\checkmark$ 98.6 14.53+5.3
+Table 2: Comparison of our CaTFormer against other end-to-end methods on the Brain4Cars dataset, using interior and exterior streams, with F1-score (%) and parameters (M).
 Implementation Details
-Our proposed CaSTFormer was implemented by PyTorch and experiments were performed on a server with six NVIDIA RTX 2080 Ti GPUs.
+Our proposed CaTFormer was implemented by PyTorch, and experiments were performed on a server with six NVIDIA RTX 2080 Ti GPUs.
 The model was trained end‐to‐end on Brain4Cars using the Adam optimizer (initial learning rate $1\times 10^{-3}$) for 160 epochs with a batch size of 16.
-During training, each input comprised a chunk of frames randomly sampled from the 5-second pre-maneuver segment; at inference, chunks were obtained via uniform sampling.
+During training, each input comprised a chunk of frames randomly sampled from the 5-second pre-maneuver segment.
+When testing, chunks were obtained via uniform sampling.
 The weight $\alpha$ in the unified loss was empirically set to 0.1.
@@
 In driving intention prediction, straight driving is considered background, and only turns and lane changes are treated as target events.
-To evaluate our CaSTFormer model, we define the following metrics based on the predictions: true positives (TP: correctly predicted maneuvers), false positives (FP: maneuvers predicted incorrectly as another maneuver), false positive predictions (FPP: predicting a maneuver when none occurred), and missing predictions (MP: failing to detect an actual maneuver).
+To evaluate our CaTFormer model, we define the following prediction-based metrics: true positives (TP: correctly predicted maneuvers), false positives (FP: maneuvers misclassified as another maneuver), false optimistic predictions (FPP: predicting a maneuver when none occurred), and missing predictions (MP: failing to detect an actual maneuver).
 Given the set of all behaviors $\mathcal{G}$ and target maneuvers $\mathcal{G}^{\prime}=\mathcal{G}\setminus\{\mathrm{straight}\}$, Precision (Pr), Recall (Re), and F1-score are computed as follows:
-$\mathrm{Pr}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+FP_{m}+FPP_{m}},$ (13)
-$\mathrm{Re}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}^{\prime}}\frac{TP_{m}}{TP_{m}+MP_{m}},\quad F_{1}=\frac{2*\mathrm{Pr}*\mathrm{Re}}{\mathrm{Pr}+\mathrm{Re}}.$
-Comparison with State-of-the-art Methods
-Table 2: Comparison of our CaSTFormer against other end-to-end methods on the Brain4Cars dataset, using internal and external streams, with F1-score (%) and parameters (M).
-Method Inside Outside F1-score (%) Param. (M)
-Gebert (Gebert et al. 2019) $\checkmark$ 81.7 85.3+162
-$\checkmark$ 43.4 85.3+162
-$\checkmark$ $\checkmark$ 73.2 170.5+162
-Rong (Rong, Akata, and Kasneci 2020) $\checkmark$ 75.5 46.2+162
-$\checkmark$ 66.4 5.4+162
-$\checkmark$ $\checkmark$ 84.3 57.9+162
-TIFN (Guo et al. 2023) $\checkmark$ $\checkmark$ 87.9 12.3+5.3
-IDIPN (Liu et al. 2025) $\checkmark$ $\checkmark$ 94.5 11.75+5.3
-CaSTFormer (Ours) $\checkmark$ $\checkmark$ 98.6 14.53+5.3
-Table 1 provides a systematic comparison of both single- and multi-modal methods on the Brain4Cars dataset.
-Notably, our camera-only CaSTFormer variant achieves an F1-score of 97.6% (precision 96.7%, recall 98.5%), markedly surpassing all previous single-modality methods such as DCNN (92.1%) and CF-LSTM (92.1%).
-When enriched with speed information, CaSTFormer attains a new state-of-the-art F1-score of 98.6% (precision 98.7%, recall 98.5%), outperforming the best prior multi-modal model, FedPRM (95.2% F1), by 3.4%.
-These results demonstrate that CaSTFormer not only establishes a new performance standard but does so with fewer sensor inputs, highlighting its efficiency and robustness in driving intention prediction.
-Figure 4 presents the confusion matrices of our CaSTFormer and TIFN (Guo et al.
-2023).
-CaSTFormer yields a sharper diagonal and substantially fewer off-diagonal entries, demonstrating its superior discrimination of similar maneuvers and reduced false predictions.
-Table 3: F1-scores on the Brain4Cars dataset for evaluation on video segments truncated 1–4 s prior to action onset.
+$\mathrm{Pr}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}}\frac{TP_{m}}{TP_{m}+FP_{m}+FPP_{m}},$ (13)
+$\mathrm{Re}$ $=\frac{1}{|\mathcal{G}^{\prime}|}\sum_{m\in\mathcal{G}}\frac{TP_{m}}{TP_{m}+MP_{m}},\ F_{1}=\frac{2*\mathrm{Pr}*\mathrm{Re}}{\mathrm{Pr}+\mathrm{Re}}.$
 Method F1-score (%)
 [-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
-Rong (Rong, Akata, and Kasneci 2020) (in) 75.7 73.1 68.6 58.5 48.2
-Rong (Rong, Akata, and Kasneci 2020) (out) 66.4 62.4 47.0 38.8 38.9
-Rong (Rong, Akata, and Kasneci 2020) (both) 84.3 78.9 70.6 60.3 53.4
+Rong (Rong et al. 2020) (in) 75.7 73.1 68.6 58.5 48.2
+Rong (Rong et al. 2020) (out) 66.4 62.4 47.0 38.8 38.9
+Rong (Rong et al. 2020) (both) 84.3 78.9 70.6 60.3 53.4
 TIFN (Guo et al. 2023) 87.9 80.9 71.0 55.0 44.6
 IDIPN (Liu et al. 2025) 94.5 84.1 74.2 62.0 55.4
-CaSTFormer (Ours) 98.6 97.4 90.1 78.4 63.7
+CaTFormer (Ours) 98.6 97.4 90.1 78.4 63.7
+Table 3: F1-scores on the Brain4Cars dataset for evaluation on video segments truncated 1–4 s before action onset.
+Comparison with State-of-the-art Methods
+Table 1 provides a systematic comparison of both single- and multi-modal methods on the Brain4Cars dataset.
+Notably, our camera-only CaTFormer variant achieves an F1-score of 97.6% (precision 96.7%, recall 98.5%), markedly surpassing all previous single-modality methods such as DCNN (92.1%) and CF-LSTM (92.1%).
+When enriched with speed information, CaTFormer attains a new state-of-the-art F1-score of 98.6% (precision 98.7%, recall 98.5%), outperforming the best prior multi-modal model, FedPRM (95.2% F1), by 3.4%.
+These results demonstrate that CaTFormer not only establishes a new performance standard but does so with fewer sensor inputs, highlighting its efficiency and robustness in driving intention prediction.
+Fig.
+4 presents the confusion matrices of our CaTFormer and TIFN (Guo et al.
+2023).
+CaTFormer yields a sharper diagonal and substantially fewer off-diagonal entries, demonstrating its superior discrimination of similar maneuvers and reduced false predictions.
 Detailed comparative results between our method and other end-to-end approaches on full-video inputs appear in Table 2.
@@
 This result reflects the intrinsic trade-off between early intervention and predictive accuracy.
-Our CaSTFormer consistently achieves superior performance across all truncated settings, demonstrating its robustness in driving intention prediction.
+Our CaTFormer consistently achieves superior performance across all truncated settings, demonstrating its robustness in driving intention prediction.
 Figure 4: The confusion matrix tested on Brain4cars dataset. Left is ours, right is the result of TIFN (Guo et al. 2023). The color deepens as the value increases.
-Ablation Study
-Effect of components in CaSTFormer.
-To evaluate the contribution of each component in CaSTFormer, we conducted systematic ablation studies with results presented in Table 4.
-Using a dual-stream Transformer baseline (T-Base) that independently encodes internal and external features through simple concatenation, we progressively introduced the RSF, CPE and FSN modules to assess their individual impact on F1-score performance.
-We further explored various combined configurations to understand module interactions and their cumulative effects.
-The experiments reveal consistent performance improvements across all configurations, validating that each module makes meaningful contributions to driving intention prediction while their synergistic combination yields additive enhancement benefits.
-Table 4: F1-scores on the Brain4Cars dataset between the dual-stream Transformer baseline and progressively augmented variants that introduce RSF, CPE and FSN.
 Model F1‐score (%)
 [-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
-T-Base 95.8 94.2 85.4 73.7 63.2
-T-Base+RSF 97.1 95.6 87.1 75.2 61.9
-T-Base+CPE 97.0 95.4 86.9 74.9 61.1
-T-Base+FSN 96.6 94.9 86.3 73.9 62.6
-T-Base+RSF+CPE 97.4 95.7 87.4 75.9 60.3
-T-Base+RSF+FSN 98.0 96.7 88.5 76.8 65.6
-T-Base+CPE+FSN 97.8 96.4 88.0 76.2 62.4
-T-Base+RSF+CPE +FSN(CaSTFormer) 98.6 97.4 90.1 78.4 63.7
-Figure 5: F1‐scores (%) of different attention designs on the Brain4Cars dataset.
-Effect of attention design in CaSTFormer.
-To evaluate the effect of different attention mechanisms on driving intention prediction, we conducted ablation experiments on the Brain4Cars dataset comparing three attention schemes, as summarized in Table 5.
+Base 95.8 94.2 85.4 73.7 63.2
+Base+R 97.1 95.6 87.1 75.2 61.9
+Base+C 97.0 95.4 86.9 74.9 61.1
+Base+F 96.6 94.9 86.3 73.9 62.6
+Base+R+C 97.4 95.7 87.4 75.9 60.3
+Base+R+F 98.0 96.7 88.5 76.8 65.6
+Base+C+F 97.8 96.4 88.0 76.2 62.4
+CaTFormer (R+C+F) 98.6 97.4 90.1 78.4 63.7
+Table 4: F1‐scores on the Brain4Cars dataset for the dual-stream Transformer baseline (Base) and its variants augmented with RDF (R), CRE (C), and FSN (F).
+[…]
+Result Visualization
+Fig.
+5 demonstrates that the model employs a temporal attention mechanism to realize a full reasoning path from dynamic event understanding to static decision attribution.
+Temporally, (a) and (b) delineate broad, task‑relevant event windows (e.g., the lane‑change interval), whereas (c) and (d) concentrate on a small set of decisive frames, accentuating discriminative cues and pinpointing instantaneous triggers.
+This sequence closely mirrors human cognition in which one perceives an event in its entirety before pinpointing its core cause.
+In the spatial dimension, (e) and (f) anchor the model’s reasoning in semantically relevant regions, including the driver’s facial state inside the cabin and the key road environment outside, entirely consistent with real‑world logic.
+This process vividly demonstrates how the model integrates critical cues to arrive at a reliable judgment, thereby substantiating the soundness of its decision‑making.
+Ablation Study
+Effect of components in CaTFormer.
+To evaluate the contribution of each component in CaTFormer, we conduct systematic ablation studies, as shown in Table 4.
+Starting from a dual-stream Transformer baseline (Base) that performs late fusion via feature concatenation, we progressively add RDF, CRE, and FSN to measure their impact on F1-score.
+We further explored various combinations to explore module interactions and their cumulative effects.
+Experimental results confirm that each module improves intention prediction, with additional gains from their integration.
+Figure 6: F1-scores (%) of various attention mechanisms on the Brain4Cars dataset, offset by 15 units for visual clarity.
+Effect of attention design in CaTFormer.
+To evaluate the effect of different attention mechanisms on driving intention prediction, we conducted ablation experiments on the Brain4Cars dataset comparing three attention schemes, as summarized in Fig.
+6.
 The results indicate that our proposed Bidirectional Dependency Attention (BDA) more effectively captures spatio-temporal correlations between interior and exterior streams while suppressing noise, thereby demonstrating its superiority in isolating dynamic cues and enhancing overall model robustness.
-Table 5: F1‐scores (%) of different the sequential correspondence orders in CaSTFormer on the Brain4Cars dataset.
-Module Sequence F1‐score (%)
-[-5,0] [-5,-1] [-5,-2] [-5,-3] [-5,-4]
-[…]
-[…]
-[…]
-[…]
-Effect of module sequencing in CaSTFormer.
-[…]
 Figure 7: F1‐scores (%) for different values of $\alpha$ in the loss function on the Brain4Cars dataset.
 Effect of intention loss.
-To assess the influence of the intention‑loss weight $\alpha$ in our composite objective, we conducted an ablation study by varying $\alpha$ to identify the optimal balance between core spatio‑temporal feature learning and dedicated intention supervision.
-The results in Fig.
-6 show that setting $\alpha=0.10$ produces the highest F1‑scores across most anticipation horizons, because this value injects enough gradient from the intention loss to enrich the discriminative power of spatio‑temporal embeddings with causal cues while preserving the stability and generalization capacity of the primary classification objective.
-Smaller values of $\alpha$ under‑utilize intention supervision and fail to capture these pre‑action signals, whereas larger values allow the intention loss to dominate optimization, creating gradient conflicts that degrade both spatio‑temporal coherence and overall anticipation accuracy.
+We study the effect of the intention-loss weight $\alpha$ through an ablation analysis.
+As shown in Fig.
+7, $\alpha=0.10$ achieves the best F1-score, balancing temporal learning with intention supervision and providing informative gradients without destabilizing the main objective.
+Smaller $\alpha$ under-supervises subtle pre-action cues, whereas larger $\alpha$ induces gradient conflicts that degrade temporal coherence and anticipation accuracy.
 Conclusion
-In this paper, we propose CaSTFormer, an interpretable prototype-driven causal spatio-temporal transformer for driving intention prediction.
-Our approach extracts dual‐stream interior and exterior features and processes them through a structured pipeline to align and fuse multi‐modal representations.
-Extensive experiments on the Brain4Cars dataset demonstrate that CaSTFormer achieves state‐of‐the‐art prediction accuracy with only a marginal increase in model size, validating its efficiency and suitability for real‐time driver assistance.
+In this paper, we introduced CaTFormer, a unified architecture that explicitly models causal interactions between driver behavior and environmental context for accurate intention prediction.
+Our approach extracts dual-stream features and merges them through a structured fusion pipeline.
+Extensive experiments on the Brain4Cars dataset confirm that CaTFormer achieves state-of-the-art accuracy, demonstrating its suitability for real-time driver assistance.
+Acknowledgments
+This work was supported in part by the Beijing Jiaotong University Research Fund under Grant KKA309004533.